OpenStudy (anonymous):

Find c such that the function : f(x)=x^2−6 xc is continuous everywhere. C=?

OpenStudy (abb0t):

Set them equal to each other. There's a theorem that explains this but i 4got the name.

OpenStudy (anonymous):

\[f(x)=\begin{cases}x^2-6 &\text{for $x<c$}\\ 8x-22& \text{for $x>c$}\end{cases}\] I'm sure you mean ≤ or ≥ for one of those pieces, so I'll assume you meant \[f(x)=\begin{cases}x^2-6 &\text{for $x\le c$}\\ 8x-22& \text{for $x>c$}\end{cases}\] For f(x) to be continuous at c, the following must be satisfied: \[\lim_{x\to c^-}f(x)=\lim_{x\to c^+}f(x)\] Do you know where to take it from here?